OpenStudy (anonymous):

The function H described by H(x) = 2.75x+71.48 can be used to predict the height in centimeters of a woman whose humerus (the bone from the elbow to the shoulder) is x cm long. predict the height of a woman whose humerus is 34 cm long.

OpenStudy (amistre64):

i made a monkey :) \begin{array}c&&*&*&*&*&&&&&&&&&&&&&&&\\&*&&&&&*&&&&&&&&&&&&&&\\*&&*&&*&&*&&&&&&&&&&&&&&\\*&*&&&&*&*&&&&&&&&&&&&&&\\*&&*&*&*&&*&&&&&&&&&&&&&&\\&*&&&&*&&&&&&&&&&&&&&&\\&&*&*&*&&*&*&&&&&&&&&*&*&*&&\\&&&&&*&&&*&&&&&&&*&&&&*&\\&*&*&&&*&&&*&&&&&&&*&&&&*&\\*&&*&&&&*&&&*&&&&*&*&&&&*&&\\*&&*&&&&*&&&&&&*&&&&&*&&&\\&&&*&*&*&&&&&&&*&&&&&*&&*&\\&&&*&&&&&&&&*&&&&&&&*&&\\&&&*&&&&&&&&*&&&&&&&&&\\&&*&&*&*&&&&&&*&&&&&&&&&\\&*&&&&&*&&&&*&&&&&&&&&&\\&*&&&&&*&&&&*&&&&&&&&&&\\&*&&&&&*&&&*&&&&&&&&&&&\\&&*&&&*&&&*&&&&&&&&&&&&\\*&*&*&&&*&*&*&&&&&&&&&&&&&\\\end{array}

OpenStudy (anonymous):

if H(x) = 2.75x + 71.48 represents the height in cm of a woman given that x is the length of the humerus, just plug 34cm into H for x H(34) = 2.75(34) + 71.48 = 164.98cm hope that helps

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no problem